# Topology of Fermi seas and geometry of their boundaries for free particles in one and two-dimensional lattices

### v1

@@
Free gases of spinless fermions moving on a geometric background with lattice symmetries are considered. Their Fermi seas and corresponding boundaries may be classified according to their topological properties at zero temperature. This is accomplished by considering the flat orbifolds ${\mathbb{R}}^{d}/\Gamma$, with $\Gamma$ being the crystallographic group of symmetry in $d$-dimensional momentum space. For $d=1$, there are 2 topological classes: a circumference, corresponding to an insulator and an interval, identified as a conductor. For $d=2$, the number of topological classes extends to 17: there are 8 with the topology of a disk identified as conductors and 4 corresponding to a 2-sphere matching insulators, both sets eventually including finite numbers of conical singularities and reflection corners at the boundaries. The rest of the listing includes single cases corresponding to insulators (2-torus, real projective plane, Klein bottle) and conductors (annulus, Möbius strip). Physical interpretations of the singularities are provided, as well as examples that fit within this listing. Given the topological nature of this classification, its results are expected to be robust against small perturbative interactions.
@@ 3 The $2d$ case
This hamiltonian does not have rotational invariance in momentum space, but rather the discrete symmetries $p_{x}\leftrightarrow p_{y}$, $p_{x}\leftrightarrow-p_{x}$ and $p_{y}\leftrightarrow-p_{y}$. Now the Fermi sea has the shape of a diamond in momentum space. The first BZ is given by the square of side intervals $(-\pi/a,\pi/a)$ in both the $p_{x}$ and $p_{y}$ axes. The Fermi momentum $p_{F}$ is given by half the length of the side of the diamond. These diamond shapes have four corners at an angle of $\pi/2$ that may be regarded as Van Hove singularities, such that the Fermi velocity (the gradient of the energy with respect to the momentum) vanishes there [8]. This is the generic picture for $\nu<1$, which describe conductors generically. For example, for $\nu=1/2$, the vertices of the diamond are located at $(\pm\pi/a,\pm\pi/a)$. Thus, the geometry of the Fermi sea for $\nu<1$ is that of a diamond, and generically describes a conductor. Its topology is that of a disk, with singular boundary. As in the $d=1$ case, a special configuration exists for $\nu=1$. The Fermi sea extends beyond the first BZ, with vertices of the diamond located at $\sqrt{2}(-\pi/s,\pi/a)$ and $p_{F}=\pi/a$. This means that the opposite sides of the Fermi sea can be mapped one into another by the addition of a vector proportional to $(\pm 2\pi/a,\pm 2\pi/a)$, meaning that they are equivalent in the lattice and, therefore, identified (the so called nesting mechanism). The topology of the Fermi sea becomes that of a $2d$ torus $T^{2}$, with no boundary and interpreted as an insulator. This is just one example out of the many that may be explicitly considered in $2d$.
@@
The last three columns of Table 2 contain the topological classes of the Fermi seas and their boundaries. A schematic representation of every orbifold listed in Table 2 is provided in [17] in pages 79-80. There are 7 basic topologies: the 2-sphere ($S^{2}$), the 2-disk ($D^{2}$), the real projective plane $\mathbb{R}P^{2}$, the 2-torus ($T^{2}$), the Klein bottle ($K$), the annulus ($A$) and the Möbius strip ($M$). All of these are Euclidean planar spaces, such that their Euler characteristics is zero (flat orbifolds are known as parabolic). This property is automatic in the last 4 classes but not in the first 3. Indeed, the real projective plane needs a couple of conical singularities of order 2 to become flat. Moreover, the spaces $S^{2}$ and $D^{2}$ also require a prescribed number of isolated singularities so that the resulting space is flat: this results in 4 $S^{2}$ and 8 $D^{2}$ subclasses. These possibilities are determined from the general formula of the Euler characteristic for an orbifold $O$, which is given by (see, e.g., [15]):

### v2

@@
Free gases of spinless fermions moving on a geometric background with lattice symmetries are considered. Their Fermi seas and corresponding boundaries may be classified according to their topological properties at zero temperature. This is accomplished by considering the flat orbifolds ${\mathbb{R}}^{d}/\Gamma$, with $\Gamma$ being the crystallographic group of symmetry in $d$-dimensional momentum space. For $d=1$, there are 2 topological classes: a circumference, corresponding to an insulator and an interval, identified as a conductor. For $d=2$, the number of topological classes extends to 17: there are 8 with the topology of a disk generically identified as conductors and 4 corresponding to a 2-sphere matching insulators, both sets eventually including finite numbers of conical singularities and reflection corners at the boundaries. The rest of the listing includes single cases corresponding to insulators (2-torus, real projective plane, Klein bottle) and conductors (annulus, Möbius strip). Physical interpretations of the singularities are provided, as well as examples that fit within this listing. Given the topological nature of this classification, its results are expected to be robust against small perturbative interactions.
@@ 3 The $2d$ case
This hamiltonian does not have rotational invariance in momentum space, but rather the discrete symmetries $p_{x}\leftrightarrow p_{y}$, $p_{x}\leftrightarrow-p_{x}$ and $p_{y}\leftrightarrow-p_{y}$. For generic $\nu<1$ the Fermi sea has the shape of a rounded diamond in momentum space, but for $\nu=1/2$, it has the shape of a proper diamond. The first BZ is given by the square of side intervals $(-\pi/a,\pi/a)$ in both the $p_{x}$ and $p_{y}$ axes. For $\nu=1/2$ the Fermi momentum $p_{F}$ is given by half the length of the side of the diamond. It has four corners at an angle of $\pi/2$, located at $(\pm\pi/a,\pm\pi/a)$, that may be regarded as Van Hove singularities, such that the Fermi velocity (the gradient of the energy with respect to the momentum) vanishes there [8]. Thus, the geometry of the Fermi sea for $\nu<1$ is that of a rounded or sharp diamond, and generically describes a conductor. As in the $d=1$ case, a special configuration exists for $\nu=1$. The Fermi sea extends beyond the first BZ, with vertices of the diamond located at $\sqrt{2}(-\pi/s,\pi/a)$ and $p_{F}=\pi/a$. This means that the opposite sides of the Fermi sea can be mapped one into another by the addition of a vector proportional to $(\pm 2\pi/a,\pm 2\pi/a)$, meaning that they are equivalent in the lattice and, therefore, identified (the so called nesting mechanism). The topology of the Fermi sea becomes that of a $2d$ torus $T^{2}$, with no boundary and interpreted as an insulator. This is just one example out of the many that may be explicitly considered in $2d$.
@@
The last three columns of Table 2 contain the topological classes of the Fermi seas and their boundaries. A schematic representation of every orbifold listed in Table 2 is provided in [17] in pages 79-80. There are 7 basic topologies: the 2-sphere ($S^{2}$), the 2-disk ($D^{2}$), the real projective plane $\mathbb{R}P^{2}$, the 2-torus ($T^{2}$), the Klein bottle ($K$), the annulus ($A$) and the Möbius strip ($M$). All of these are Euclidean planar spaces, such that their Euler characteristics is zero (flat orbifolds are known as parabolic). This property is automatic in the last 4 classes but not in the first 3. Indeed, the real projective plane needs a couple of conical singularities of order 2 to become flat. Moreover, the spaces $S^{2}$ and $D^{2}$ also require a prescribed number of isolated singularities so that the resulting space is flat: this results in 4 $S^{2}$ and 8 $D^{2}$ subclasses. These possibilities are determined from the Riemann–Hurwitz formula of the Euler characteristic for an orbifold $O$, which is given by (see, e.g., [15]):